OpenStudy (anonymous):

What is the sum of the first five terms of a geometric series with a1 = 6 and r = 1/3? i got 242/81 is that correct?

OpenStudy (cwrw238):

= 6 * (1.3)^5 - 1 1/243 - 1 ---------- = -------- - 2/3 1/3 - 1 = - 242 -3 ---- * -- 243 2

OpenStudy (cwrw238):

oh - that * 6

OpenStudy (anonymous):

= - 242 -3 ---- * -- so I find out what this is and that should be my answer? 243 2

OpenStudy (anonymous):

= - 242 -6 ---- * -- did u mean -6 instead of -3? 243 2

OpenStudy (cwrw238):